### v1

# A Class of Countably Covered Two-Dimensional Sofic Shifts

Abstract
A multidimensional sofic shift is called countably covered if it has an SFT cover containing only countably many configurations. In contrast to the one-dimensional setting, not all countable sofic shifts are countably covered. We study a subclass of countable shift spaces and characterize the countably covered sofic shifts among them.
1 Introduction
A multidimensional shift of finite type (SFT) is a set of colorings of $\mathbb{Z}^{d}$ using finitely many colors that is defined by a finite number of local rules. A sofic shift is the image of an SFT, called its cover, under a map that identifies some of the colors with each other. We study the following problem: given a two-dimensional sofic shift whose cardinality is countable, when does it admit an SFT cover that is also countable?
We concentrate on the case where $d=2$ and the sofic shift is vertically constant, that is, each coordinate of a vertical column must have the same color. This is the setting in which the author previously proved in [10] that not all countable sofic shifts have countable SFT covers. The author proved a necessary condition and showed it to be sufficient it in a restricted subclass of sofic shifts. In this article we prove the same statement in another subclass: those in which all but finitely many columns have the same “zero” color, and the number of nonzero columns is restricted by a computable function of the size of the smallest gap between two non-null columns.
It is a well known result of Aubrun-Sablik [1] and Durand-Romashchenko-Shen [4] that the $\mathbb{Z}^{2}$-lift of a $\mathbb{Z}$-shift space is sofic, i.e. has an SFT cover, if and only if it is effectively closed. In [3], Durand and Romashchenko further prove that if the $\mathbb{Z}$-shift space is quasiperiodic, then the SFT cover can be taken quasiperiodic as well, and the same holds for minimality. One can say that quasiperiodicity and minimality are properties that can always be lifted to the SFT cover of a vertically constant sofic shift. Countability, as it turns out, cannot always be lifted in this way.
On the other hand, this study is also related to the equal entropy problem of whether every $\mathbb{Z}^{2}$-sofic shift has an SFT cover of equal topological entropy. Desai showed in [2] that there always exist covers whose entropy is arbitrarily close to that of the sofic shift, but reaching the exact value remains an open question. In particular, it is not known whether every zero entropy sofic shift admits a zero entropy SFT cover, that is, whether having zero entropy is a liftable property. Coutability, like having zero entropy, is a “smallness property” of shift spaces, whose liftability we investigate.
This article is dedicated to prof. Jarkko Kari, my former PhD advisor, on his 60th birthday. I am thankful to him for many years of collaboration and friendship.
2 Preliminaries
2.1 Definitions
We recall some basic notions of multidimensional symbolic dynamics. Let $A$ be a finite alphabet and $d\geq 1$ a dimension. The $d$-dimensional full shift over $A$ is the set $A^{\mathbb{Z}^{d}}$, whose element are called configurations. It is equipped with the shift action of $\mathbb{Z}^{d}$, defined by $\sigma^{\vec{v}}(x)_{\vec{n}}=x_{\vec{n}+\vec{v}}$ for $x\in A^{\mathbb{Z}^{d}}$ and $\vec{v},\vec{n}\in\mathbb{Z}^{d}$. A $d$-dimensional pattern with finite domain $D\subset\mathbb{Z}^{d}$ is a map $P:D\to A$. Each pattern defines a cylinder set $[P]=\{x\in A^{\mathbb{Z}^{d}}\mid x|_{D}=P\}$, and these sets form a basis for the profinite topology. A topologically closed and shift-invariant subset of $A^{\mathbb{Z}^{d}}$ is called a shift space. If $X\subseteq Y$ are shift spaces, we say $X$ is a subshift of $Y$.
Each shift space $X\subseteq A^{\mathbb{Z}^{d}}$ is defined by a set $F$ of forbidden patterns, in the sense that $X=\bigcap_{P\in F}\bigcap_{\vec{v}\in\mathbb{Z}^{d}}\sigma^{v}(A^{\mathbb{Z}^{d}}\setminus[P])$. If $F$ can be chosen finite, then $X$ is a shift of finite type, or SFT; if $F$ can be chosen computably enumerable, then $X$ is effectively closed. If $X$ is a shift space and $\pi:A\to B$ is any map to another alphabet $B$, the coordinate-wise projection $\pi(X)\subseteq B^{\mathbb{Z}^{d}}$ is also a shift space. If $X$ is an SFT, then $\pi(X)$ is a sofic shift, and the pair $(X,\pi)$ is its SFT cover. It is well known that all sofic shifts are effectively closed, and that the converse does not hold when $d\geq 2$.
A configuration $x\in A^{\mathbb{Z}^{2}}$ is vertically constant if $\sigma^{(0,1)}(x)=x$. A shift space is vertically constant if all its configurations are. The (two-dimensional) lift of a one-dimensional configuration $x\in A^{\mathbb{Z}}$ is the vertically constant configuration $x^{\dagger}\in A^{\mathbb{Z}^{2}}$ with $(x^{\dagger})_{(i,j)}=x_{i}$ for all $(i,j)\in\mathbb{Z}^{2}$. The lift of a shift space $X\subseteq A^{\mathbb{Z}}$ is $X^{\dagger}=\{x^{\dagger}\mid x\in X\}$.
Theorem 1 ([4, 1]).
For $d\geq 2$, every effectively closed vertically constant $\mathbb{Z}^{d}$-shift space is sofic.
A one-dimensional configuration $x\in A^{\mathbb{Z}}$ is ultimately periodic to the right, if there exists $p>0$ such that $x_{i}=x_{i+p}$ holds for all large enough $i>0$. Ultimate periodicity to the left is defined symmetrically.
A function $f:\mathbb{N}\to\mathbb{N}$ is lower semicomputable if there exists a computable function $g:\mathbb{N}^{2}\to\mathbb{N}$ with $g(n,k+1)\geq g(n,k)$ for all $n,k\in\mathbb{N}$ and $f(n)=\max_{k\in\mathbb{N}}g(n,k)$ for all $n\in\mathbb{N}$. The function $g$ is called a lower approximation of $f$. Upper semicomputability is defined analogously.
2.2 Previous work
We recall the countable cover conditions from [10].
Definition 2.
Let $X\subseteq A^{\mathbb{Z}}$ be a one-dimensional shift space. we say $X$ satisfies the countable cover conditions if
  1. 1.
    $X$ is effectively closed,
  2. 2.
    every $x\in X$ is ultimately periodic to the left and right, and
  3. 3.
    given $u,v,w\in A^{+}$, it is decidable whether ${}^{\infty}uvw^{\infty}\in X$.
Theorem 3 ([10]).
Let $X\subset A^{\mathbb{Z}}$ be a one-dimensional shift space. If the lift $X^{\dagger}\subset A^{\mathbb{Z}^{2}}$ is a countably covered sofic shift, then $X$ satisfies the countable cover conditions.
In [10] it was shown that the converse holds if $X$ is a subshift of a one-dimensional countable sofic shift. Note that one-dimensional countable sofic shifts have a very constrained structure: their configurations consist of periodic regions of bounded periods, with a bounded number of period-breaking patterns in between [7, Lemma 4.8].
3 Gap width shifts
We define a class of countable one-dimensional shift spaces. For an alphabet $A$ with $0\in A$, denote $A_{0}=A\setminus\{0\}$. For a word $w\in A^{*}$, denote by $|w|_{\neq 0}$ the total number of nonzero symbols in $w$.
Definition 4.
Let $A$ be a finite alphabet containing 00, and $f:\mathbb{N}\to\mathbb{N}$ any function. The gap width shift $G(A,f)\subset\{0,1\}^{\mathbb{Z}}$ is the shift space defined by the forbidden patterns
$F(f)=\left\{w=ua0^{n}bv\in\{0,1\}^{*}\;\middle|\;n\geq 0,a,b\in A_{0},|w|_{\neq 0}>f(n)\right\}.$
If $A$ is clear from the context, we may denote $G(A,f)=G(f)$.
The gap width shift $G(f)$ contains exactly those configurations $x\in A^{\mathbb{Z}}$ for which the number of nonzero symbols is at most $f(n)$, where $n$ is the minimal length of a finite run of 0-symbols in $x$. In other words, the width of the smallest gap between nonzero symbols puts a bound on the number of gaps. It follows that $G(f)$ only contains configurations with a finite (but usually unbounded) number of nonzero symbols.
We are interested in subshifts of gap width shifts, and whether their lifts are sofic or countably covered. For the full gap width shifts $G(f)$, soficness of the lift corresponds to a computability-theoretic property of the function $f$.
Proposition 5.
The lift $G(f)^{\dagger}$ is sofic if and only if $f:\mathbb{N}\to\mathbb{N}$ is upper semicomputable.
Proof.
By Theorem 1, $G(f)^{\dagger}$ is sofic if and only if $G(f)$ is effectively closed.
Suppose $f$ is upper semicomputable with computable approximation $g:\mathbb{N}^{2}\to\mathbb{N}$. Then $F(f)$ is computably enumerable, since the condition $|w|_{\neq 0}>f(n)$ is equivalent to $\exists k\in\mathbb{N}:|w|_{\neq 0}>g(n,k)$. Hence $G(f)$ is effectively closed.
Suppose then that $G(f)$ is effectively closed. Denote $F=\{0,1\}^{*}\setminus\mathcal{L}(G(f))$, which is a computably enumerable set. For each $n\in\mathbb{N}$ we have $f(n)=\min\left\{k\in\mathbb{N}\;\middle|\;a(0^{n}a)^{k}\in F\right\}$, where $a\in A_{0}$ is arbitrary. It follows that $f$ is upper semicomputable. ∎
As our main result, we show that the converse of Theorem 3 holds for subshifts of gap width shifts with computable gap functions. The proof is deferred to Section 5.
Theorem 6.
Let $f:\mathbb{N}\to\mathbb{N}$ be a computable function and $X\subseteq G(f)$ a shift space. Then $X^{\dagger}$ is a countably covered sofic shift if and only if $X$ satisfies the countable cover conditions.
Corollary 7.
The lift $G(f)^{\dagger}$ is a countably covered sofic shift if and only if $f:\mathbb{N}\to\mathbb{N}$ is computable.
Proof.
Suppose that $G(f)^{\dagger}$ is countably covered, so that $G(f)$ satisfies the countable cover conditions. Now $f(n)=\min\left\{k\in\mathbb{N}\;\middle|\;{}^{\infty}0a(0^{n}a)^{k}0^{\infty}\notin G(f)\right\}$ for each $n\in\mathbb{N}$, where $a\in A_{0}$ is arbitrary. Since ${}^{\infty}uvw^{\infty}\in G(f)$ is decidable by assumption, $f$ is computable.
Suppose then that $f$ is computable. Then $G(f)$ satisfies the countable cover conditions: it is effectively closed since the set of forbidden patterns $F(f)$ is computably enumerable, every configuration is eventually all-00 in both directions, and ${}^{\infty}uvw^{\infty}$ can be easily decided by measuring the smallest gap $110^{n}$ 1, computing $f(n)$ and comparing that to the number of $1$-symbols. ∎
In particular, the subclass of gap width shifts with countably covered sofic lifts is also characterized by a computability condition, and is neither empty nor equal to the class of gap width shifts with sofic lifts.
Definition 8.
Let $X\subseteq G(f)$ be a shift space. Define the gap function of $X$ by $f_{X}(n)=\max\{|x|_{\neq 0}\mid x\in X,m\leq n,a0^{m}b\sqsubset x\}$ for $n\geq 0$.
In other words, the gap function $f_{X}:\mathbb{N}\to\mathbb{N}$ is the pointwise smallest nondecreasing function such that $X\subseteq G(f_{X})$.
Proposition 9.
Let $f:\mathbb{N}\to\mathbb{N}$ be any nondecreasing function and suppose $X\subseteq G(f)$ satisfies the countable cover conditions. Then $f_{X}$ is lower semicomputable.
Proof.
As $X$ satisfies the countable cover conditions, we can algorithmically enumerate its configurations in the form ${}^{\infty}0v0^{\infty}$ for $v\in A^{*}$. Every such $v$ containing an occurrence of $a0^{n}b$ is a witness for $f_{X}(n)\geq|v|_{\neq 0}$, and $f_{X}(n)=|v|_{\neq 0}$ holds for at least one $v$. ∎
Example 10.
There exists a subshift of a gap width shift that satisfies the countable cover conditions, but whose gap function is not computable. Theorem 6 does not apply in this situation; we do not know whether such shift spaces have countably covered lifts.
Define a shift space $X\subset\{0,1\}^{\mathbb{Z}}$ as the set of all translates of the following configurations: the all-00 configuration; configurations ${}^{\infty}0(10^{n})^{k}10^{\infty}$ for all $n\geq 2$ and $0\leq k\leq n$; and configurations ${}^{\infty}010^{n}(10^{k})^{k}10^{\infty}$ for all $n\geq 2$ such that the $n$th Turing machine halts on empty input in exactly $k>n$ steps. This set is a shift space. Since $f_{X}$ has Busy Beaver-like growth, it is not computable.
We claim that $X$ satisfies the countable cover conditions. First, it is effectively closed: every word with at most two $1$-symbols occurs in $X$; a word of the form $0^{a}(10^{n})^{k}0^{b}$ with $k\geq 2$ occurs in $X$ if and only if $2\leq n$ and $k\leq n$; and all other words that occur in $X$ have the form $0^{a}10^{n}(0^{k}1)^{m}0^{b}$ for some $a,b\geq 0$ and $2\leq n<k$ and $2\leq m\leq k$, and their occurrence can be checked by simulating the $n$th Turing machine for $k$ steps. Second, every configuration of $X$ is clearly ultimately periodic. Third, whether ${}^{\infty}uvw^{\infty}\in X$ can be decided by checking that it is of the correct form, and possibly simulating a Turing machine for a bounded number of steps.
4 Construction toolbox
In this section we present several auxiliary constructions of countable SFTs that are used as part of the main construction. Many of them were used in some form in [10], but we restate them here both for completeness and for ease of reuse.
Our most fundamental building block is the grid shift, a standard example of a nontrivial two-dimensional SFT. In our constructions it acts as a way of transferring information from one part of a configuration to everywhere else, as a scaffold on which we can implement computation, and as a “clock” that periodically resets said computation.
Example 11 (Grid shift).
The grid shift $X_{\mathrm{grid}}\subset A_{\mathrm{grid}}^{\mathbb{Z}^{2}}$ is the SFT defined on the Wang tile set $A_{\mathrm{grid}}$ shown in Figure 1. A configuration of $X_{\mathrm{grid}}$ either contains a regular grid of horizontal and vertical lines, or contains no parallel lines. In particular, $X_{\mathrm{grid}}$ is a countable SFT. For each $n\geq 1$, let $x_{\mathrm{grid}}(n)\in X_{\mathrm{grid}}$ be the configuration containing a grid of side length $n$ with a crossing of grid lines at the origin.
Figure 1: A sample configuration of the grid shift $X_{\mathrm{grid}}$. Dotted lines are tile borders, along which colors must match.
The following is essentially [10, Example 14]. It allows us to constrain the number of nonzero columns in a configuration of $(A^{\mathbb{Z}})^{\dagger}$ by the size of a grid on a separate layer.
Lemma 12 (Counting columns).
Fix a finite alphabet $A$ containing 00. There exists an SFT $X_{\mathrm{count}}\subset(A^{\mathbb{Z}})^{\dagger}\times X_{\mathrm{grid}}\times A_{\mathrm{count}}^{\mathbb{Z}^{2}}$ with the following properties.
  1. 1.
    For each $n\in\mathbb{N}$, the set $X_{\mathrm{count}}(n):=\left\{z\in X_{\mathrm{count}}\;\middle|\;\pi_{\mathrm{grid}}(z)=x_{\mathrm{grid}}(n)\right\}$ is countable, and the projection $\pi_{A}(X_{\mathrm{count}}(n))$ is exactly the set of configurations of $(A^{\mathbb{Z}})^{\dagger}$ with at most $n-2$ nonzero columns.
  2. 2.
    For each $x\in(A^{\mathbb{Z}})^{\dagger}$ the set $\left\{z\in X_{\mathrm{count}}\;\middle|\;\pi_{A}(z)=x\right\}$ is countable.
Note that we could enforce a bound of $n$ nonzero columns in item 1 instead of $n-2$, but this would require a slightly more complex tile set.
Proof.
The alphabet $A_{\mathrm{count}}$ consists of two layers of Wang tiles, which are shown in Figure 2. The first layer is the counter layer, and the second one is the synchronization layer. The alphabet of $X_{\mathrm{count}}$ also contains an $A$-layer and a grid layer. Each layer must respect the local matching rules of its Wang tiles. In addition, the layers are constrained by the following rules, where $c\in A\times A_{\mathrm{grid}}\times A_{\mathrm{count}}$ is an arbitrary symbol:
  1. 1.
    The grid layer of $c$ contains a horizontal line if and only if its counting layer contains a solid horizontal line (a tile from group 3 in Figure 2). In this case, the counting layer contains a white dot if and only if the synchronization layer contains a dashed vertical line.
  2. 2.
    The grid layer of $c$ contains a vertical line if and only if the synchronization layer contains a solid vertical line.
  3. 3.
    If the counting layer of $c$ contains a solid vertical line (a tile from group 2 in Figure 2), then the grid layer must contain a vertical line.
  4. 4.
    The grid layer of $c$ contains both a vertical and a horizontal line if and only if the counting layer is in both groups 2 and 3.
  5. 5.
    If the counting layer of $c$ has a dashed horizontal line on its west border (so the tile is in group 1 of Figure 2), then the line turns north if and only if the $A$-layer is nonzero.
We describe the structure of configurations $z\in X_{\mathrm{count}}$; see Figure 3 for reference. Suppose that the grid layer of $z$ contains a grid of finite side length $n$, that is, $z\in X_{\mathrm{count}}(n)$. By rules 1 and 2, this grid completely determines the positions of all solid horizontal lines of the counting layer and all solid vertical lines of the synchronization layer. By rule 4 it also determines the positions of solid vertical lines of the counting layer, which must coincide with those of the other layers. The bi-infinite stripe between each pair of solid horizontal lines of the counting layer contains a single bi-infinite dashed path. By rule 5 the path goes up by one tile whenever it meets a nonzero column on the $A$-layer, and continues horizontally otherwise. Since there are $n-1$ rows between two adjacent grid lines, the $A$-layer of $z$ can contain at most $n-2$ nonzero columns.
The positions of the dashed paths of different stripes in $z$ are synchronized by the diagonal lines of the counting layer: when the path crosses a solid vertical line, a diagonal line is sent to the southeast, which eventually meets a solid horizontal line and produces a white dot. By rule 1 the positions of the dots must match the intersections of the dashed lines of the synchronization layer and the horizontal lines of the grid layer. Thus there are a finite number of options for the positions of the dashed paths, each of which determines the synchronization layer. Since there are also finitely many choices for the $A$-layer, the set $X_{\mathrm{count}}(n)$ is countable. We also see that the $A$-layer can be chosen freely, as long as it has at most $n-2$ nonzero columns. This proves item 1.
For item 2, fix a configuration $x\in(A^{\mathbb{Z}})^{\dagger}$, and let $z\in X_{\mathrm{count}}$ be such that $\pi_{A}(z)=x$. We show that there are a countable number of choices for $z$. If the grid layer of $z$ contains a grid of finite side length, the claim follows from item 1 Suppose it does not contain such a grid. Then the counting layer contains at most one solid horizontal line and at most one solid vertical line, and the synchronization layer contains at most one solid vertical line. In particular, there are countably many choices for the contents of the synchronization layer. Since there must be a solid horizontal line between each pair of dashed paths on the counting layer, $z$ contains at most two such paths. There are a countable number of choices for the number of these paths and the positions at which they cross the y-axis. Since we fixed the $A$-layer $\pi_{A}(z)=x$, after these choices the dashed paths are completely determined. Finally, between two diagonal lines of the counting layer there must be a thick horizontal line, so the number of choices for them is likewise countable. ∎
(a)123(b)
Figure 2: The two layers of the alphabet $A_{\mathrm{count}}$ of the SFT $X_{\mathrm{count}}$: (a) the counting layer and (b) the synchronization layer.
[…]
Figure 3: A sample configuration of $X_{\mathrm{count}}$. The counting layer is shown, as are the dashed lines of the synchronization layer. The two arrows mark nonzero columns of the $A$-layer.
Example 13 (Counter machines).
We describe a way of simulating computation by nondeterministic oracle counter machines in SFTs. We do not explicitly define a $k$-counter machine, but it consists of a finite oracle alphabet $A$, a finite state set $Q$, and initial state $q_{0}\in Q$, and a transition relation $\delta$. The machine has $k$ counters that store one natural number each, which are initialized to 0, and possibly an infinite oracle tape $x\in A^{\mathbb{N}}$. If the oracle tape exists, then one of the counters is designated as the oracle counter. Alternatively, the oracle tape may be a bi-directional configuration $x\in A^{\mathbb{Z}}$, in which case each counter will store an integer instead of a natural number. In this latter case the counter machine is called bidirectional.
In one computation step the machine will:
  • check the relative order of the counters, and which of them have value 0,
  • check the symbol $x_{n}$, where $n$ is the current value of the oracle counter,
  • based on the above data, the current state, and possibly a nondeterministic choice, change the value of any counters by at most 1 and assume a new state.
The machine may also have a designated final state $q_{f}$. If the machine enters $q_{f}$, it accepts, and if it ever has no possible next state, it rejects.
Let $M$ be a one-directional $k$-counter machine as above. Define $H_{M}=H^{\prime}_{M}\times\{{\leftarrow},{\rightarrow}\}$ and $A_{M}=\{\#\}\cup((H_{Q}\cup\{L,R\})\times\{Z,P\}^{k}\times A)$, where $H^{\prime}_{M}$ is an auxiliary alphabet. Define an SFT $X_{M}\subset A_{M}^{\mathbb{Z}^{2}}$ as follows. If a horizontal row of $x\in X$ contains a $-\ell$ segment of non-##\## symbols starting at $\vec{v}$, then a length-$(\ell-1)$ segment must start below it at $\vec{v}-(0,1)$ and a length-$(\ell+1)$ segment above it at $\vec{v}+(0,1)$. We interpret time as increasing upward and introduce (nondeterministic) rules that transform a segment into the one above it.
Each segment consists of $k+1$ layers. Its first layer must be in the language $L^{*}H_{M}R^{*}$ (storing the position, direction and current transition in a “zig-zag head”) and the next $k$ layers must be in $P^{*}Z^{*}$ (storing the value of one counter). The final layer is in $A^{*}$, and each segment is a prefix of the one above; their limit corresponds to the oracle tape. The head travels along the segment in the direction of its arrow component, with speed 2 to the east and speed 1 to the west. When it steps east over a counter, it updates the counter’s value based on its $H^{\prime}_{M}$-component; when it steps west over one, it records its place in the ordering of the counters, as well as the symbol on the $A$-layer in case of the oracle counter. When it reaches the east border of the segment, it turns west, and upon reaching the west border, it updates its $H^{\prime}_{M}$-component based on the information it stores. In this way, one west-east sweep corresponds to one computation step of $M$.
If the machine is bidirectional, we modify the computation cone to grow both to the left and to the right. The left and right halves of the cone have different background colors, and the central column of the cone corresponds to counter value 00.
The most important property of $X_{M}$ is that if we ignore the $A$-component, the set of “degenerate” configurations – those that do not contain a simulated computation of $M$ – is countable. Indeed, such a configuration has at most one connected region of non-##\## cells, which is shaped like a half-plane and contains at most one infinite sweep of the zig-zag head. We will modify this basic construction as needed, in particular by allowing the computation cones to be cut short when $M$ halts, and attaching them to grid cells.
Recall from [6, Section 11.2] that any Turing machine computation can be simulated by a counter machine computation in a way that incurs an exponential blowup in the number of computation steps. In our construction, the first $n$ simulated steps of a counter machine computation fit in an exponential-size square pattern. Hence, the blowup from a Turing machine computation to a pattern is doubly exponential.
Example 14 (Simulating tile sets in grids).
Let $(T_{n})_{n\in\mathbb{N}}$ be a computable sequence of Wang tile sets whose colors are in $\mathbb{N}$, meaning that each $T_{n}$ is a finite subset of $\mathbb{N}^{4}$. We decribe a way of simulating each $T_{n}$ by a single SFT, which is countable if each $T_{n}$ is.
Take the grid shift $X_{\mathrm{grid}}$ and superimpose on each gray region a computation of a counter machine using a variation of the SFT $X_{M}$ in which the computation cones are allowed to be cut off at any point after the simulated machine has halted, and the oracle tape is blank. We give $M$ one counter that corresponds to the value of $n$, and four counters that correspond to the four cardinal directions. The machine $M$ first nondeterministically sets each of these counters to some value (by incrementing them some number of times) and never modifies them again. Then it simply checks whether the 5-tuple $(n,t_{e},t_{n},t_{w},t_{s})$ they form satisfies $(t_{e},t_{n},t_{w},t_{s})\in T_{n}$, halting if it does.
We also use signals to transfer information between the machines in adjacent grid cells: we enforce each simulated machine in a single configuration to choose the same value of $n$, and for the direction counters to be related as in a Wang tiling (so the north counter of a grid cell must equal the south counter of its north neighbor, and analogously for west and east). See [9, Theorem 6.2] for a more detailed construction and analysis.
Example 15 (Coupling simulated tile sets with vertical layers).
We describe a modification to the SFT $X$ of Example 14 that adds a vertically constant layer over a separate alphabet $A$ and gives the simulated tile sets access to it. For this, we assume that each tile set $T_{n}$ has a separate layer over the auxiliary alphabet $B=\{0,1,2\}\cup(A\times\{0,1\})$, and all $2\times 1$ patterns except $00$, $0a$, $a1$, $a2$, $11$, $12$ and $22$ for $a\in A\times\{0,1\}$ are forbidden on this layer. This layer is used to “probe” the contents of the $A$-layer.
We construct the counter machine $M$ so that it contains a special probe state $q_{b}$ for each $b\in B$. It will enter a probe state $q_{b}$ at some point during its computation if and only if the $B$-layer of the tile it simulates equals $b$.
We superimpose on $X\times(A^{\mathbb{Z}})^{\dagger}$ a configuration over the tile set $A_{\mathrm{probe}}$ of Figure 4, called the probe layer, and further constrain it using the following rules:
  1. 1.
    The thick horizontal and vertical lines of the probe layer must coincide with those of the grid layer.
  2. 2.
    The highlighted tiles in Figure 4, which include only those with $b^{\prime}\in A\times\{0\}$, are called the probe tiles. The $A$-part of the symbol $b^{\prime}$, $a_{0}$ or $a_{1}$ of a probe tile must be equal to the symbol on the $A$-layer.
  3. 3.
    If the west border of a grid cell has on its right the zig-zag head of a simulated machine $M$ in a probe state $q_{b}$ for $b\in B$, then the corresponding vertical line of the probe layer must hold the symbol $b$.
Denote by $H\subseteq X\times(A^{\mathbb{Z}})^{\dagger}\times A_{\mathrm{probe}}^{\mathbb{Z}^{2}}$ the SFT thus obtained. The probe layer of a sample configuration is depicted in Figure 5.
Consider a configuration $(x,y,z)\in H$, where $x\in X$ contains a grid of width $m\geq 0$ that simulates a configuration of tile set $T_{n}$ for some $n\in\mathbb{N}$. By item 1, the horizontal and vertical lines of the probe layer $z$ are aligned with the grid, and by item 3 and our assumption on $M$, each vertical line segment in $z$ holds the value of the $B$-layer of the tile of $T_{n}$ simulated on its right. By our assumption on $T_{n}$, each row of the simulated tiling contains at most one symbol from $A\times\{0,1\}$, and on its right some number of $1$-symbols, all other symbols coming from $\{0,2\}$. Thus, each row of the grid contains at most one vertical segment $L$ holding a value $(a,c)\in A\times\{0,1\}$ on the probe layer.
If the next $k>0$ segments to the right hold $1$-symbols, the rightmost one of these emits a dashed line to the left on the southmost row of the grid cell. This signal is raised to the north by one step each time it crosses a segment holding a $1$-symbol, and upon collision with $L$, sends a diagonal signal to the southeast. This signal carries the same symbol $(a,c)$, and as it hits the south border of the grid cell at distance $k$ from segment $L$, rule 2 guarantees that the value of the $A$-layer at that position is $a$. We also see that $k<m$. Because of the black disks at the borders of certain tiles, $c=1$ if and only if the diagonal signal hits the tile neighboring the next segment to the right from $L$, if and only if $k=m-1$. In the case $k=0$, by rule 2 the segment $L$ holds the symbol $(a,0)$, where $a\in A$ is the symbol of the $A$-layer on the same column. In this way, the $B$-layer of the simulated tilings can be used to access the width of the grid layer and the full contents of the $A$-layer.
As for the cardinality of the new SFT, note first that if the grid layer contains a finite-width grid, then the probe layer is completely determined by the other layers. Conversely, if there is no finite-width grid, then the probe layer too contains no parallel thick lines, and a simple case analysis shows that the number of valid contents for the probe layer is countable when the other layers are fixed.
[…]
Figure 4: The tile set $A_{\mathrm{probe}}$. Here $b$ ranges over $B$, $b^{\prime}$ ranges over $\{0,1,2\}\cup(A\times\{0\})$, $a$ ranges over $A$, and $a_{i}$ for $i\in\{0,1\}$ ranges over $A\times\{i\}$.
[…]
Figure 5: A configuration over $A_{\mathrm{probe}}$, superimposed over a configuration of $X_{\mathrm{grid}}$. Each vertical segment is labeled with the type of symbol it carries; $a_{i}$ stands for $A\times\{i\}$.
A weaker version of the following result was stated explicitly in [10].
Lemma 16.
Let $g:\mathbb{N}\to\mathbb{N}$ be a partial computable function. Then there exists a sequence $(C_{n},T_{n},a_{n},b_{n},t_{n})_{n\in\mathbb{N}}$, computable in polynomial time, with the following properties for all $n\in\mathbb{N}$.
  1. 1.
    $T_{n}$ is an NW-deterministic Wang tileset with edge colors $C_{n}=\{1,\ldots,k_{n}\}$, where $k_{n}=O(\log\log n)$.
  2. 2.
    $a_{n},b_{n}\in C_{n}$ are edge colors and $t_{n}\in T_{n}$ a tile such that for some $m\geq g(n)$, $T_{n}$ tiles an $m\times m$ rectangle whose entire north border is colored with $a_{n}$, west border with $b_{n}$, and the tile at the southeast corner is $t_{n}$. Furthermore, $T_{n}$ does not tile any such rectangle for $m<g(n)$.
  3. 3.
    $T_{n}$ does not tile the infinite plane.
Proof.
For $n\in\mathbb{N}$, let $M$ be a Turing machine that counts to some number $k\geq g(n)$, enters a special state $q$ that is not used otherwise, counts to $k$ again and then halts. We construct the machine in such a way that it has $\lceil\log n\rceil+K$ states for some constant $K$, by having it first write $n$ in binary on its tape and then perform a uniform algorithm with input $n$.
In [5], Kari constructs a NW-deterministic version of the aperiodic Robinson tile set [8] and embeds an arbitrary Turing machine in it. Its tilings consist of nested square patterns, called $(2^{n}-1)$-squares, each of which hosts a simulation of the embedded machine for approximately $n/2$ steps. If the machine halts in these $n$ steps, the pattern cannot be tiled correctly.
Denote by $R_{n}$ this tile set with $M_{n}$ being the embedded machine. It can tile a rectangle of side length $g(n)$ whose entire north border is colored by one color $a_{n}$ – the one that occurs at the north border of every southeast-facing $(2^{n}-1)$-square – and west border by the color $b_{n}$ found on the west border of these squares. The tile that simulates the special state $q$ is chosen as $t_{n}$; we can build $M_{n}$ and the simulation so that it always lies on the diagonal of a $(2^{n}-1)$-square. Then item 2 is satisfied. Since the machine $M_{n}$ eventually halts, $R_{n}$ does not tile the plane. Finally, the number of colors in $R_{n}$ grows linearly in the number of states and transitions of $M_{n}$, and we can make the latter grow as $O(\log\log n)$ by replacing $g$ with $g\circ\exp$ and using $M_{\lceil\log n\rceil}$ in place of $M_{n}$. Note that the sequence is allowed to have repeated values. ∎
The following lemma is a key ingredient of our construction. It allows us to take a small grid and use it to enforce the existence of a large but still finite grid on another layer.
Lemma 17 (Blowing up grids).
Let $g:\mathbb{N}\to\mathbb{N}$ be a total computable function. There exists a finite alphabet $A_{g}$ and a countable SFT $X_{g}\subset X_{\mathrm{grid}}\times X_{\mathrm{grid}}\times A_{g}^{\mathbb{Z}^{2}}$ such that for each $n\geq 1$, there exists a nonempty finite set $K\subset\mathbb{N}\cap[g(n),\infty)$ such that
$\left\{y\in X_{\mathrm{grid}}\;\middle|\;\exists z:(x_{\mathrm{grid}}(n),y,z)\in X_{g}\right\}=\left\{\sigma^{\vec{v}}(x_{\mathrm{grid}}(k))\;\middle|\;k\in K,\vec{v}\in n\mathbb{Z}^{2}\right\},$
and if $(x,y,z)\in X_{g}$ is such that $y$ has a grid of finite width, then so does $x$.
Proof.
Let $(C_{n},T_{n},a_{n},b_{n},t_{n})_{n\in\mathbb{N}}$ be the sequence of tile sets and tiles given by Lemma 16 for the function $n\mapsto g(2^{n})$. Construct a new sequence of tile sets $(S_{n})_{n\in\mathbb{N}}$ as follows. Take the grid shift $X_{\mathrm{grid}}$, and decorate each tile without a horizontal or vertical line with a tile of $T_{n}$ in all possible ways. Two adjacent decorated tiles must respect the adjacency rules of $T_{n}$. If the north neighbor of a decorated tile $t\in S_{n}$ is a horizontal line, then the north color of the $T_{n}$-decoration of $t$ must be $a_{n}$, and if the west neighbor of $t$ is a vertical line, then the west color of the decoration must be $b_{n}$. If the east neighbor of $t$ is a vertical line and its south neighbor is a horizontal line, then its decoration must be $t_{n}$. The colors of all other edges of the $T_{n}$-decorations are not constrained. This concludes the definition of $S_{n}$.
The tile set $S_{n}$ does not admit tilings without a finite grid, since $T_{n}$ does not tile the infinite plane. By compactness, the set of grid sizes it admits is finite. It does admit at least one tiling with a grid of some width $m\geq g(2^{n})$, but not of any width $m<g(2^{n})$. Also, the decorations inside each grid cell forms the same pattern in any tiling, since $T_{n}$ is NW-deterministic and the colors north and west borders of the grid cells are fixed. Thus the set of valid tilings over $S_{n}$ is countable.
We describe the SFT $X_{g}$. Let $M$ be a counter machine with six named counters $b,m,t_{e},t_{n},t_{w},t_{s}$ that checks whether the quadruple $t=(t_{e},t_{n},t_{w},t_{s})\in\mathbb{N}^{4}$ is a tile of $S_{m}$, and that $b=1$ when $t$ corresponds to a tile with a horizontal grid line and $b=0$ otherwise. Let $h(m)\geq m$ be an upper bound for the number of computation steps required for this check for a fixed $m$ and any values of the other named counters. By Lemma 16 and the exponential slowdown in simulating a Turing machine by a counter machine, we can choose $h(m)=2^{m^{q}}$ for some $q\in\mathbb{N}$. We may also assume that the machine always takes the same number of steps for the check when $m$ is fixed. For example, we may have it perform the check for all possible values of the counters $t_{e}$, $t_{n}$, $t_{w}$ and $t_{s}$ that are below the maximum value $\log\log m$ and ignore all results except the one with the actual counter values; the total number of steps is still $2^{n^{O(1)}}$, so this does not violate the previous assumption. After $M$ has finished the check, it runs for at least $h(m+1)$ more steps (e.g. by simply counting to $h(m+1)$) and then halts.
Similarly to Example 14, let $X\subset X_{\mathrm{grid}}\times X_{\mathrm{grid}}\times A_{g}^{\mathbb{Z}^{2}}$ be the SFT where each grid cell of the first component contains a simulated computation of $M$ on the third component, and adjacent grid cells synchronize the simulated counters so that each machine has the same value of $m$ and the colors of the simulated tiles match. We allow the simulation of $M$ to be cut at any point after the check but before it has halted. For each cross tile $t$ of the first component, we require that $t$ is matched with a horizontal line on the second component if and only if $t$ is adjacent to a simulated $b$-counter with value $1$ on the $A_{g}$-component. Also, each cross tile of the second component must be matched with a cross tile on the first component. For finitely many $n$ (including those for which the grid cells of $x_{\mathrm{grid}}(n)$ are too small to host a computation), we can enforce the contents of the second component by explicit rules. This concludes the definition of $X_{g}$. We call the three components of $X_{g}$ the small grid layer, the large grid layer and the simulation layer.
We show that $X_{g}$ has the desired properties. Take any $n\geq 1$. Let $y$ and $z$ be such that $z^{\prime}=(x_{\mathrm{grid}}(n),y,z)\in X_{g}$. If $n$ is large enough, every grid cell of the small grid hosts a simulation of $M$ that computes a tile from the same set $S_{m}$ and enforces the adjacency rules to its neighbors. We must have $m\geq\log n$, since $M$ always runs for at least $m$ steps, which corresponds to at least $2^{m}$ rows in the tiling.
By Lemma 16, every valid tiling over $S_{m}$ contains a grid of finite width. Hence the small grid layer contains regularly spaces cross tiles that must be matched with horizontal lines on the large grid layer $y$. Thus $y$ has a finite-width grid, that is, it is a shifted version of some $x_{\mathrm{grid}}(k)$ for $k\in\mathbb{N}$. Moreover, the simulated grid of $S_{m}$ has width at least $g(2^{m})\geq g(2^{\log n})=g(n)$, so that of $y$ has width at least $ng(n)>g(n)$ for large enough $n$ (and small values of $n$ are handled separately). As the two grids must be aligned, we have $y=\sigma^{\vec{v}}(x_{\mathrm{grid}}(k))$ for some $k\geq g(n)$ and $\vec{v}\in n\mathbb{Z}^{2}$. Since $x_{\mathrm{grid}}(n)$ is $n$-periodic both horizontally and vertically, $\vec{v}$ can be chosen arbitrarily. Thus there exists some set $K\subset\{g(n),g(n+1),\ldots\}$ such that the claimed equation holds.
By compactness, the set $K$ must be finite. We show that for large enough $n$ it is also nonempty. For a given $m$, the simulated machine takes at most $h(m)$ steps to perform its computations, after which it runs for at least $h(m+1)$ more steps before halting, where $h(m)=2^{m^{q}}$. It takes $C^{t}$ rows to simulate $t$ computation steps of the machine for a constant $C\geq 2$. This means that it is enough to find an $m$ such that $C^{h(m)}\leq n\leq C^{h(m+1)}$, so that the machine has enough space to perform the check in an $n\times n$ square, but not enough to halt. This is equivalent to
$m\leq\left(\log\left(\log_{C}n\right)\right)^{1/q}\leq m+1.$
Clearly, such an $m$ can always be found. Thus there exist $y$ and $z$ with $(x_{\mathrm{grid}}(n),y,z)\in X_{g}$, or in other words, $K\neq\emptyset$.
Finally, suppose $(x,y,z)\in X_{g}$ is such that $y$ contains a finite-width grid. Since every cross of $y$ is matched with a cross in $x$, the latter must also contain a finite-width grid. ∎
5 Proof of the main result
We now combine our puzzle pieces in order to form a proof of Theorem 6.
Proof of Theorem 6.
One direction follows from Theorem 3. For the converse, suppose we are given a nondecreasing computable function $f:\mathbb{N}\to\mathbb{N}$ and a shift space $X\subseteq G(f)$ that satisfies the countable cover conditions. Our goal is to construct a countable SFT cover for $X^{\dagger}$. In the first phase of the construction, we embed $X^{\dagger}$ into a countably covered sofic shift $Y$ that is contained in $G(g)^{\dagger}$ for some function $g:\mathbb{N}\to\mathbb{N}$ that majors $f$. After this, we will further restrict the SFT cover of $Y$, transforming it into a cover of $X$.
We begin to construct a countable SFT $Z$ and a projection $\pi:Z\to A^{\mathbb{Z}^{2}}$ such that $Y=\pi(Z)$ satisfies $G(f)^{\dagger}\subseteq Y\subseteq G(g)$ for some function $g$. This SFT consists of several layers, the first of which is the vertically constant $A$-layer $(A^{\dagger})^{\mathbb{Z}}$. We define $\pi$ as the projection to this layer.
The second layer is a copy of $X_{\mathrm{grid}}$, which we call the small grid layer. The third layer is the width restriction layer given by the tiles visible in Figure 6 and their obvious matching rules. Its thick horizontal lines must coincide with those of the small grid layer, and its vertical lines must coincide with the nonzero columns of the $A$-layer. It restricts the side length of the small grid layer to be at most the maximum distance between two nonzero columns of the $A$-layer. Note that the width restriction layer is completely determined by the previous layers in the case that there are at least two nonzero columns in the $A$-layer, and has countably many configurations otherwise.
Figure 6: A sample configuration of the width restriction layer.
Next, we use Lemma 17 to add another copy of $X_{\mathrm{grid}}$, called the large grid layer, and couple its width to that of the small grid layer using the function $f+2$. If the small grid layer has a grid of some finite width $n\in\mathbb{N}$, now the large grid layer must have a grid of finite width $m\geq f(n)+2$. We choose $g(n)$ as the maximum value of $m$, the existence of which is guaranteed by Lemma 17.
Finally, we use Lemma 12 to add a counting layer that couples the large grid layer with the $A$-layer. It has the effect of constraining the number of nonzero columns on the $A$-layer to be at most $m-2$, but places no other restrictions on it.
By the construction and the lemmas used, $Z$ is a countable SFT. It remains to be shown that $G(f)^{\dagger}\subseteq Y\subseteq G(g)^{\dagger}$. For the first inclusion, let $y\in G(f)^{\dagger}$. If $y$ contains at most one nonzero column, then the other layers of $Z$ can be filled by degenerate configurations. If it contains two or more nonzero columns, let $n>0$ be the minimum distance between any two of them. We can use $x_{\mathrm{grid}}(n)$ as the small grid layer, and then the width restriction layer can be filled legally. By Lemma 17, we can use $x_{\mathrm{grid}}(m)$ for some $m\geq f(n)+2$ as the large grid layer, and since $y\in G(f)^{\dagger}$, this does not violate the coupling given by Lemma 12. Hence $y\in Y$.
Let then $y\in Y$ be arbitrary, and let $z\in Z$ be such that $y=\pi(z)$. If the small grid layer of $z$ does not contain a finite-width grid, then $y$ contains at most one nonzero column, so $y\in G(g)^{\dagger}$. Suppose then that the small grid layer contains a grid of width $n$. Then Lemma 17 guarantees that the large grid layer contains a grid of width $m\leq g(n)$, and Lemma 12 states that there are at most $m-2$ nonzero columns in $y$. The width restriction layer ensures that the distance between any two such columns cannot be less than $n$. Hence $y\in G(g)^{\dagger}$.
We now begin the second phase of the construction. This involves adding new layers to the SFT cover $Z$ to produce a new countable SFT $Z^{\prime}$ with $\pi(Z^{\prime})=X^{\dagger}$. We first add a new copy of $X_{\mathrm{grid}}$ called the simulation grid, on which we simulate a sequence of tile sets $(T_{n})_{n\in\mathbb{N}}$ as in Example 15; they will have access to the $A$-layer and are responsible for checking its correctness. Here, $n$ is the width of the small grid layer. The size of the simulation grid is at least $t(n)\cdot m$ for a computable function $t:\mathbb{N}\to\mathbb{N}$ growing fast enough that the counter machines always have enough time to correctly simulate tiles of $T_{n}$ (recall that $m$ is the size of the large grid); we enforce this using Lemma 17. We require that the small grid, large grid and simulation grid all be aligned, so that the counter machines running on the simulation grid cells can access $n$ and the contents of the counting layer.
We define the simulated tile sets $T_{n}$ for $n\in\mathbb{N}$ as SFTs, which can then be recoded into sets of Wang tiles. The set $T_{n}$ contains $k_{n}=\binom{f(n)}{2}$ Wang tile layers, one for each two-element subset $\{a,b\}\subset\{0,1,\ldots,f(n)-1\}$; we identify these numbers with the nonzero columns of the $A$-layer. Exactly one of these layers is active at a time, and the others are inactive. For a pair of adjacent horizontal rows over $T_{n}$, the active layer must follow the adjacency rules of its layer, while the remaining $k_{n}-1$ layers must be identical across the two rows. The role of the active layer is passed on cyclically, so that on a height-$k_{n}$ horizontal strip of tiles, each layer is active on ecaxtly one row. Hence, each layer is “stretched” by a factor of $k_{n}$.
Each of the $k_{n}$ layers consists of several sublayers. The first sublayer, called the computation sublayer, contains a simulation of a deterministic bidirectional counter machine $C_{n}$. Its purpose is to repeatedly probe the $A$-layer of $Z$ and determine whether it is in $X^{\dagger}$. The second sublayer, called the anchor sublayer, is a vertically constant layer over the alphabet $\{0,1,2\}$, in which the horizontal patterns $10$, $20$ and $21$ are forbidden. The positions of the $1$-symbols on this layer are controlled by the counter machine that performs the simulation of $T_{n}$. The idea is that the $1$-symbols on the anchor sublayer of the layer corresponding to the pair $\{a,b\}$ span the region delimited by the $a$th and $b$th nonzero columns of the $A$-layer, so that the counter machine of the computation sublayer can be initialized near the $a$th column and check the configuration between it and the $b$th column.
In addition to these sublayers, the $k_{n}$ layers share a common sublayer over the alphabet $B$ of Example 15, which is used to probe the $A$-layer. The counter machine of each layer has two special counters $c_{1}$ and $c_{2}$. The $1$-symbols of the common $B$-layer must always coincide with the portion of the computation cone that is between the ends of these counters on the active layer. When the zig-zag head of the active layer steps on the $A\times\{0,1\}$-symbol of the $B$-layer, it can store it as part of its state.
The machine $C_{n}$ behaves as follows. First, it determines the number $m$ of $1$-symbols on the anchor sublayer. Then, using the $B$-layer via its special counters $c_{1}$ and $c_{2}$, it probes for the width $h(n)$ of the simulation grid and then the full contents of the $A$-layer on the union of all the grid cells that simulate $1$-symbols on the anchor layer. In this way it receives a word $w\in A^{m\cdot h(n)}$. It checks whether $|w|_{\neq 0}\leq f(d)$, where $d$ is the minimum distance between two nonzero symbols in $w$, rejecting (and producing a tiling error) if this is not the case. Next, $C_{n}$ checks whether ${}^{\infty}0w0^{\infty}\in X$, which is possible since $X$ satisfies the countable cover conditions. If this holds, then $C_{n}$ halts and accepts. If ${}^{\infty}0w0^{\infty}\notin X$, then by compactness there exists a number $p\geq 0$ such that $0^{p}w0^{p}\notin\mathcal{L}(X)$, which we can also compute since $X$ is effectively closed. After this, the machine probes the $p$ symbols to the left and right of $w$ on the $A$-layer. If the result is $0^{p}w0^{p}$, then $C_{n}$ rejects; otherwise, it accepts.
We now describe how the counter machines of the simulation grid determine the positions of the $1$-symbols of the anchor layer. The simulation grid is aligned with the small and large grids, and the counter machine in each cell of the simulation grid has access to the contents of them and the counting layer. In particular, it can read the size $n$ of the small grid, as well as the leftmost position of a dashed vertical line on the synchronization layer in its grid cell. Denote by $e$ the distance of this line from the west border of the grid cell. The machine checks that $e<f(n)$, rejecting if not. Then, on each layer $\{a,b\}$, on the anchor sublayer it places a 00-symbol if $e<a$, a $1$-symbol if $a\leq e\leq b$, and a $2$-symbol if $b<e$. In this way we guarantee that every configuration contains at most $f(n)$ nonzero columns, and that each pair of these columns has a corresponding simulated layer on which the space between them is highlighted with $1$-symbols.
We add two more sublayers to each of the $k_{n}$ layers of $T_{n}$, the purpose of which is to periodically restart the computation of the machine $C_{n}$ on the computation layer. The third sublayer is a grid, which is coupled with the run of $1$-symbols on the second sublayer similarly to the width restriction layer: if $201^{m}$ 2 occurs in the second sublayer, the third sublayer must have a grid of side length at most $m$. The fourth sublayer is a grid that is coupled to the third sublayer using Lemma 17 with a computable function $h^{\prime}:\mathbb{N}\to\mathbb{N}$ such that the simulated counter machine $C_{n}$ will always take at most $h^{\prime}(m)$ rows of tiles to finish its computation if the number of $1$-symbols on the anchor layer is $m$; again, such a function clearly exists. A vertical line of the latter grid erases whatever computation was taking place on the computation sublayer, and initializes a new computation on each occurrence of $01$ on the anchor layer (of which there can be at most one).
Finally, there may be symbols $a\in A_{0}$ such that ${}^{\infty}0a0^{\infty}\notin X$. For each such symbol, there necessarily exists $p\geq 0$ such that $0^{p}a0^{p}\notin\mathcal{L}(X)$, and we forbid this horizontal pattern from the $A$-layer of $Z^{\prime}$. This concludes the construction of $Z^{\prime}$.
We first claim that $Z^{\prime}$ is countable. As $Z$ was already proved countable, it suffices to fix all the layers of $Z$ (the $A$-layer, the small and large grids, and their width restriction and counting layers) and show that the remaining layers have countably many choices. If the small grid layer contains no finite-size grid, then neither does the simulation grid layer, by the last claim on Lemma 17. Hence, it contains at most one computation cone of a counter machine, which has a countable number of possible computation paths, plus a countable number of choices for other degenerate signals. Suppose then that the small grid layer contains a grid of size $n$, so that the large grid layer contains a grid of some size $m\geq f(n)+2$. Then the simulation grid layer contains a grid of size at least $t(n)\cdot m$, on which a tiling of $T_{n}$ is simulated (since the counter machines in the grid cells can probe the value of $n$).
On each of the $k_{n}$ layers of $T_{n}$, the anchor layer is completely determined by the previously described layers. If the anchor layer does not contain a finite run of $1$-symbols, then it does not contain a finite-width grid, and neither does the fourth sublayer by Lemma 17. Hence the simulation sublayer contains at most one position where a computation of $C_{n}$ is initialized, and since $C_{n}$ is deterministic, this constitutes a countable number of choices. On the other hand, if the anchor layer contains a pattern $201^{k}$ 2, then the third sublayer contains a grid of width at most $k$, and the fourth contains a grid of width at most $h^{\prime}(m)$. The latter grid periodically restarts the computation of $C_{n}$ on the same column. Hence the number of choices is again countable.
We now prove $\pi(Z^{\prime})=X^{\dagger}$. Take any $x\in X^{\dagger}$ and use it as the $A$-layer. Since $x\in G(f)^{\dagger}$, we can safely fill all the layers of $Z$, in particular putting a grid of finite width $n$ on the small grid layer, equal to the size of the smallest gap on the $A$-layer, and one of some width $m$ on the large grid layer. We choose to fill the counting layer so that the dashed horizontal lines are directly above the horizontal grid lines on the left end of the configuration. On the simulation grid layer we can safely put a grid of finite width at least $t(n)\cdot m$ by Lemma 17. By our choice of the function $t$, the counter machines inside these grid cells have enough time to correctly simulate a tiling of $T_{n}$.
Consider a layer of $T_{n}$ corresponding to a pair $\{a,b\}$. The anchor sublayer of this layer is determined by the previous layers. If it does not contain a finite run of $1$-symbols, then the remaining sublayers can safely be filled with degenerate, uniform configurations. If it does contain a run $201^{k}$ 2, we can put a grid of width $k$ on the third sublayer, and a grid of some width at least $h^{\prime}(k)$ on the fourth.
This leaves the computation sublayer (and the common $B$-layer, which is determined by the previous layers and only constrains the computation sublayers). On this layer we are forced to initialize computations of the machine $C_{n}$ on the intersections of the horizontal grid lines of the fourth sublayer and the leftmost $1$-column on the anchor sublayer. The machine probes the small and large grid layers and the $A$-layer. Since the $A$-layer comes from $x\in X$, it cannot reject; the only way to do so would be to find too many nonzero columns or a forbidden word of the form $0^{p}w0^{p}$, neither of which can occur on the $A$-layer. Hence we have produced a valid tiling of $Z^{\prime}$.
Conversely, take any $z\in Z^{\prime}$; we claim that its $A$-layer is in $X^{\dagger}$. If the small grid layer is degenerate, then so is the large grid layer, and the $A$-layer may contain at most one nonzero column. Since we explicitly forbade any ${}^{\infty}0a0^{\infty}\notin X$ from occurring on the $A$-layer, this case is handled. Suppose hence that the small grid layer contains a grid of size $n$. Then, as in the argument for the countability of $Z^{\prime}$, we must have finite-size grids on the large grid and simulation grid layers, the latter of which must simulate a tiling of $T_{n}$, which in turn consists of the $k_{n}$ layers plus the shared $B$-layer. The simulation layer also restricts the vertical dashed lines on the counting layer of $Z$ to be at height at most $f(n)$, and assigns to each nonzero column of the $A$-layer a unique successive number in $\{0,\ldots,f(n)\}$. Let these numbers span a sub-interval $\{a,a+1,\ldots,b\}$.
We restrict our attention to the layer of the simulated $T_{n}$-tiling corresponding to the pair $\{a,b\}$. The run of $1$-symbols of its anchor sublayer is finite and spans the entire nonzero part of the $A$-layer. Thus, the third and fourth sublayers contain grids of finite size, and the computation sublayer contains an infinite number of restarted computations of $C_{n}$ that have enough time to complete by accepting. The machine $C_{n}$ reads the entire nonzero part of the $A$-layer as the word $w$. If we had ${}^{\infty}0w0^{\infty}\notin X$, then $C_{n}$ would have to find a nonzero column within a finite distance to the left or right of $w$, which is impossible, since $w$ spans all the nonzero columns. Hence we have ${}^{\infty}0w0^{\infty}\in X$, as required. ∎
6 Further results
With minor modifications to the proof of Theorem 6 we can extend it to a class of one-dimensional shift spaces that includes all subshifts of both one-dimensional sofic shifts and gap shifts. Namely, fix an alphabet $A$, a computable function $f:\mathbb{N}\to\mathbb{N}$ and a bound $b>0$. In each configuration of the gap width shift $G(f,\{0,1\})$, replace each symbol within distance $b$ from a $1$-symbol with an arbitrary symbol of $A$, and then replace each remaining (finite or infinite) run of 00-symbols with some equal-length periodic pattern over $A$ with period at most $b$. Each of these replacements is done independently of the others. Denote the set of all possible resulting configurations by $G(A,f,b)$.
We claim that subshifts of such $G(A,f,b)$ also satisfy the converse of Theorem 3. In the proof, we add a new $\{0,1\}$-layer on which we place a configuration of $G(f,\{0,1\})$ from which the $A$-layer could have been produced. This can be verified with only local rules. We use this $\{0,1\}$-layer in place of the $A$-layer in every part of the construction, except that we give the $C_{n}$-machines access to the $A$-layer in order to verify its correctness. The algorithm of $C_{n}$ is modified suitably.